OpenStudy (anonymous):

When planning his crops, Farmer Sue knows that her 15 acres can support apples and pecans. She wants to make $8,500 from her crops. She can make $1,050 per acre of apples (the variable a) and $2,500 per acre of pecans (the variable p). Which equation below would be a constraint in her system of equations? a. 1,050a + 2,500p = 8,500 b. a + p = 15 c.a − p = 15 d.2,500p − 1,050a = 8,500 HELP PLEASE <33333

OpenStudy (anonymous):

& This Oneee Please : A company sells boxes of cow bells (b) for $20 and boxes of air horns (h) for $30. They can make a batch of cow bells that fills 8 boxes or a batch of air horns that fills 9 boxes. The company only has 53 boxes. Which system of equations helps the company plan to make $275? b = 53 − h 8b + 9h = 53 8b = 53 − 9h 20b + 30h = 275 b = 53 − h 20b + 30h = 275 8b = 53 − 9h 20b − 30h = 275